OpenStudy (anonymous):

Evaluate the expression for the given value of the variable(s). (4(3h-6))/(1+h);h=-2 A. 32 B. 48 C. -48 D. 30

OpenStudy (anonymous):

\[\frac{4(3\times (-2)-6)}{1+(-2)}\] is a start

OpenStudy (anonymous):

yeah i got that it's just whenever i put my stuff in the calculator i dont get any of the answer choices.

OpenStudy (anonymous):

then \[\frac{4(-6-6)}{-1}\] \[\frac{4(-12)}{-1}\] \[\frac{-48}{-1}\] \[48\]

OpenStudy (anonymous):

don't use it, it is easier to do this with your mind than to figure out how to put it in the calculator properly, what with all the parentheses i think the point of this exercise is order of operations anyway, so calculator will weigh you down, not help

OpenStudy (anonymous):

thank you ! :)

OpenStudy (anonymous):

yw